OpenStudy (anonymous):

An astronaunt weighs 8.00x10^2 newtons on the surface of Earth. What is the weight of the astronaut 6.36x10^6 meters above the surface of Earth?

OpenStudy (anonymous):

Since g = GM/(R+r)^2 where R earth's radius which measures 6.36 * 10 ^ 6 m M is mass of the earth and G is universal gravitational constant. Also r is distance of the object from the surface of the earth. At surface of the earth r = 0 So g = GM/R^2 Now calculate g' at r = 6.36*10^2 = R weight = mg Hope this helps

OpenStudy (usman1995):

yep